one number is seven times another number. the sum of the numbers is 136. find the numbers
Found 2 solutions by checkley77, scott8148:
Answer by checkley77(12844)

x+7x=136
8x=136
x=136/8
x=17 ans. for the smaller number
7*17=119 ans. for the larger number.
Proof:
17+119=136
136=136

Answer by scott8148(6628)

n + 7n = 136

8n = 136

n = 17

7n = 119